# Divide, Interact, Sample: The Two-System Paradigm

### v1

@@ 1 Introduction
Two broad classes of MCMC algorithms are widely used: (i) Discrete-time methods based on the Metropolis-Hastings algorithm, and (ii) Continuous-time methods defined by Langevin-type stochastic differential equations (SDEs). Both approaches yield Markov chains with the desired target distribution as their stationary distribution. The increasing availability of parallel computing has made it practical to run multiple chains simultaneously, which in turn has spurred the development of ensemble-based methods that leverage interactions among chains to improve convergence.
@@
To unify these perspectives, we introduce the two-system approach. This framework divides an ensemble of $N$ particles into two interacting subsystems, with each subsystem using the empirical distribution of the other as the basis for its proposals. This symmetric coupling guarantees that the target distribution $\rho(x)$ is invariant in both the finite-$N$ setting and the infinite-particle (mean-field) limit. Starting from a mean-field Langevin SDE, we show that the two-system formulation naturally yields ensemble-chain samplers via finite-particle discretization of the continuous dynamics. By alternating updates between the two subsystems in discrete time (with appropriate Metropolis-Hastings corrections), we obtain parallel MCMC algorithms that preserve detailed balance with respect to $\rho$ and are well-suited for modern multi-core hardware.
@@ 3 Background
As mentioned above, samplers can broadly be categorized into continuous-time methods (based on Langevin-type SDEs) and discrete-time methods (e.g., the Metropolis-Hastings algorithm). Continuous-time samplers leverage the fact that a Langevin SDE has the target density $\rho(x)$ as its stationary distribution, meaning that as $t\to\infty$, the law of the SDE converges to $\rho(x)$. However, in practice, one must discretize the SDE to simulate it, and such discretizations typically introduce bias and no longer preserve $\rho(x)$ exactly. To correct this bias, the discretized SDE can be used as a proposal within a Metropolis-Hastings framework, yielding a Markov chain that converges to the correct target. More broadly, the Metropolis-Hastings algorithm provides a general recipe for constructing discrete-time samplers that converge to $\rho(x)$. In high-dimensional settings, however, naive proposals can lead to very low acceptance rates, resulting in poor mixing and inefficient exploration of the target.
@@ 3.1 Discrete-Time Samplers
The classic Metropolis-Hastings algorithm [14, 30] constructs a Markov chain $\{X_{k}\}_{k\geq 0}$ with stationary distribution $\rho(x)$. Each iteration consists of two steps: (i) Given the current state $X_{k}$, propose a new state $Y_{k}\sim Q_{C}(\cdot\mid X_{k})$, where $Q_{C}(\cdot\mid X_{k})$ is a proposal distribution parameterized by some matrix $C$ (e.g., a covariance); (ii) Accept the proposal with probability
@@ 3.2 Continuous-Time Samplers
4 A Two-System Approach for McKean-Vlasov Equations
We now develop the two-system approach for general continuous-time mean-field SDEs (McKean–Vlasov equations), of which the overdamped and underdamped Langevin equations are special cases. The key idea is to consider two coupled systems of particles such that each system “drives” the evolution of the other. In a Langevin setting, this symmetric coupling is constructed to preserve the target distribution and to directly connect the behavior of the finite-$N$ particle system with its $N\to\infty$ mean-field limit.
The mean-field Langevin equation (11) is an example of a more general McKean-Vlasov SDE, which takes the form:
@@ 4 A Two-System Approach for McKean-Vlasov Equations
with $W_{t}^{1},\ldots,W_{t}^{N}$ independent Brownian motions and $\mathbf{X}_{t}=\{X_{t}^{j}\}_{j=1}^{N}$. As $N\to\infty$, this interacting particle system converges to the mean-field equation (14) in the sense of propagation of chaos [9, 42]: the particles become asymptotically independent and each follows the law of the McKean-Vlasov solution $\overline{X}_{t}$.
@@
Under Assumptions 1 and 2, one can show that the coupled SDE (18) has a unique strong solution on $[0,T]$ (proof in Appendix A.2). Moreover, $\textrm{Law}(Z_{1,t})$ and $\textrm{Law}(Z_{2,t})$ both coincide with the original mean-field law $\mu_{t}$ for all $t\in[0,T]$. In particular, if $Z_{1,0}$ and $Z_{2,0}$ are identically distributed (with law $\mu_{0}$), then $Z_{1,t}\stackrel{{\scriptstyle d}}{{=}}Z_{2,t}$ for all $t\in[0,T]$, and this common distribution evolves according to the McKean-Vlasov equation (14) (proof in Appendix A.3). Thus, $\pi_{t}=\textrm{Law}(Z_{t})$ can be seen as the coupling of two independent copies of the mean-field solution.
@@ 5 A Two-System Approach for Samplers
The two-system framework introduces a pair of interacting Markov chains, where each chain generates proposals based on the empirical law of the other. When applied to a continuous-time mean-field sampler, this coupling preserves $\rho$ as the invariant distribution in both the finite-particle setting and the infinite-particle limit. By adding Metropolis–Hastings acceptance steps to an alternating (staggered) time-discretization of this coupled system, we obtain an ensemble-chain sampler that exactly maintains the target distribution $\rho$.
This connection demonstrates that ensemble-chain samplers approximate expectations with respect to $\rho$ by leveraging interactions between particles. In the limit as $N\to\infty$ (and as time $t\to\infty$), the ensemble statistics converge to their mean-field counterparts. This also motivates the use of adaptive samplers, which replace the whole ensemble with a single chain and rely on long-run time averages to recover the same mean-field behavior.
@@ 5.1 Continuous-Time Samplers
To illustrate, consider the mean-field overdamped Langevin example. The two-system construction leads to a coupled SDE on $Z_{t}=\{X_{t},Y_{t}\}$ (with $X_{t},Y_{t}\in\mathbb{R}^{d}$):
@@
for $t\in[0,T]$, where $\mu_{t}=\textrm{Law}(Z_{t})$ and we assume $C(\mu)$ remains bounded and positive definite for all $\mu\in\mathcal{P}(\mathbb{R}^{d})$. Under mild regularity conditions on $\nabla\log\rho$, the two-system McKean-Vlasov equation (22) admits a unique strong solution (e.g., if $\log\rho(x)$ is strictly concave outside a large ball).
@@
Then $\nabla\cdot S(Z_{t})=0$ is a sufficient condition for the corresponding Fokker-Planck to have $\rho^{\otimes 2N}$ as the invariant density [32, 34]. Since $S(Z_{t})_{i,j}$ does not depend on $(Z_{t})_{j}$, $(\nabla\cdot S(Z_{t}))_{i}=0$ for all $i\in\{1,\ldots,2N\}$. Thus $\rho^{\otimes 2N}$ is the invariant density of (23). ∎
@@ 5.2 Discrete-Time Samplers
where $\mu_{t}^{Y}$ denotes the empirical law of subsystem $Y$ at time $t$, and $\xi^{X},\xi^{Y}$ are independent standard normal increments. In other words, we first update $X$ using the current $Y$-statistics, then update $Y$ using the new $X$-statistics (this is a symmetric split step akin to a leapfrog). This yields a sequence of proposals $(X_{t},Y_{t})\to(X_{t+h},Y_{t})\to(X_{t+h},Y_{t+h})$ for the coupled system. We then apply a Metropolis–Hastings accept/reject step at each proposal to ensure detailed balance with respect to $\rho$.
@@ 6 Algorithmic Realizations of Two-System Samplers
For the ensemble-chain (multi-chain) setting, convergence to the target $\rho$ is ensured as long as each Metropolis–Hastings update satisfies detailed balance. We therefore present the two-system ensemble versions of MALA and MAKLA in Algorithms 1 and 2.
@@ 7.1 Synthetic Experiments
Coupled MAKLA delivers order-of-magnitude gains over NUTS by exploiting spatial averaging, and the adaptive variants achieve similar per-gradient efficiency, often within $10\text{\textendash}25$% of coupled MAKLA’s ESS/Grad, while using far fewer particles ($N=20$ chains versus $8d$ particles for Coupled). Under the two-system lens, this is natural: adaptive MAKLA replaces particle averaging with historical averaging. Instead of computing the preconditioner from a large, simultaneous ensemble, it estimates the same quantity from the chain’s time history. Thus, coupled and adaptive are two ways to approximate the same mean-field preconditioner; one by averaging across particles at a fixed time, the other by averaging across time with many fewer particles. Practically, this trades memory and ensemble size for a lightweight adaptation schedule, retaining most of the efficiency benefits of the coupled ensemble while substantially reducing particle count.
@@ 9 Acknowledgements
We thank Dr. Anna Lisa Varri, Dr. Amanda Lenzi, Prof. Peter Radchenko, and Deven Sethi for the helpful conversations while writing this paper. James Chok acknowledges the use of the Isaac Newton Trust (INT) grant G101121 LEAG/929. Daniel Paulin’s research was supported by an NTU Start Up Grant.
@@ A.2 Existence and uniqueness
This proof follows the standard argument of the existence and uniqueness of McKean-Vlasov SDEs, see [7, 21, 42] for more details.
@@
$|\lvert\mathbf{b}(t,Z_{t},\pi_{t})-\mathbf{b}(t,\overline{Z}_{t},\overline{\pi}_{t})\rvert^{2}\$$\leq\ 2c\Big{(}\lvert Z_{t}-\overline{Z}_{t}\rvert^{2}\ +\ \mathcal{W}_{2}^{2}(p_{1}\circ\pi_{t},p_{1}\circ\overline{\pi}_{t})+\mathcal{W}_{2}^{2}(p_{2}\circ\pi_{t},p_{2}\circ\overline{\pi}_{t})\Big{)}$ (47)
@@ A.3 Convergence in Wasserstein
If $\textrm{Law}(X_{0})=\textrm{Law}(Y_{0})=\textrm{Law}(\overline{X}_{0})$, then $\textrm{Law}(X_{t})=\textrm{Law}(Y_{t})=\textrm{Law}(\overline{X}_{t})$ for $t\in[0,T]$. That is, the mean-field dynamics of the two-system approach are the same as using one particle.
@@
$\begin{split}&dX_{t}\ =\ b(t,X_{t},\rho_{t})\,dt\ +\ \sigma(t,X_{t},\rho_{t})\,dW_{t},\\ &dY_{t}\ =\ b(t,Y_{t},\mu_{t})\,dt\ +\ \sigma(t,Y_{t},\mu_{t})\,dW_{t},\\ &d\overline{X}_{t}\ =\ b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t}\end{split}$ (66)
with $\textrm{Law}(X_{t})=\mu_{t},\,\textrm{Law}(Y_{t})=\rho_{t},\,\textrm{Law}(\overline{X}_{t})=\overline{\mu}_{t}$ and initial conditions $X_{0},Y_{0},\overline{X}_{0}$ respectively, where they share the same noise. Then for $t\in[0,T]$,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\Big{[}\mathcal{W}_{2}^{2}(\mu_{0},\overline{\mu}_{0})\ +\ \mathcal{W}_{2}^{2}(\rho_{0},\overline{\mu}_{0})\Big{]},$ (67)
@@
$\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,X_{s},\rho_{s})-\sigma(s,\overline{X}_{s}),\overline{\mu}_{s})\rvert^{2}\,ds$ (71)
@@
Since a similar inequality holds for $\mathbb{E}\lvert Y_{t}-\overline{X}_{t}\rvert^{2}$,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\$$\leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{X}_{0}\rVert^{2}\right)$ (75)
$\qquad\qquad+c\int^{t}_{0}\Big{(}\mathcal{W}_{2}^{2}(\rho_{s},\overline{\mu}_{s})\ +\ \mathcal{W}_{2}^{2}(\mu_{s},\overline{\mu}_{s})\Big{)}ds.$ (76)
Since $(X_{t})_{0\leq t\leq T}$, $(Y_{t})_{0\leq t\leq T}$, and $(\overline{X}_{t})_{0\leq t\leq T}$ are squared-integrable, it follows that $\mathbb{E}\lvert X_{t}-\overline{X}_{t}\rvert^{2}$ and $\mathbb{E}\lvert Y_{t}-\overline{X}_{t}\rvert^{2}$ are bounded for $t\in[0,T]$. Hence $\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})+\mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})$ is also bounded on $t\in[0,T]$. This allows us to apply Grönwall’s lemma
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{X}_{0}\rvert^{2}\right).$ (77)
@@
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\Big{[}\mathcal{W}_{2}^{2}(\mu_{0},\overline{\mu}_{0})\ +\ \mathcal{W}_{2}^{2}(\rho_{0},\overline{\mu}_{0})\Big{]}.$ (78)

### v2

@@ 1 Introduction
Two broad classes of MCMC algorithms are widely used: (i) Discrete-time methods based on the Metropolis--Hastings algorithm, and (ii) Continuous-time methods defined by Langevin-type stochastic differential equations (SDEs). Both approaches yield Markov chains with the desired target distribution as their stationary distribution. The increasing availability of parallel computing has made it practical to run multiple chains simultaneously, which in turn has spurred the development of ensemble-based methods that leverage interactions among chains to improve convergence.
@@
To unify these perspectives, we introduce the two-system approach. This framework divides an ensemble of $N$ particles into two interacting subsystems, with each subsystem using the empirical distribution of the other as the basis for its proposals. This symmetric coupling guarantees that the target distribution $\rho(x)$ is invariant in both the finite-$N$ setting and the infinite-particle (mean-field) limit. Starting from a mean-field Langevin SDE, we show that the two-system formulation naturally yields ensemble-chain samplers via finite-particle discretization of the continuous dynamics. By alternating updates between the two subsystems in discrete time (with appropriate Metropolis--Hastings corrections), we obtain parallel MCMC algorithms that preserve detailed balance with respect to $\rho$ and are well-suited for modern multi-core hardware.
@@ 3 Background
As mentioned above, samplers can broadly be categorized into continuous-time methods (based on Langevin-type SDEs) and discrete-time methods (e.g., the Metropolis--Hastings algorithm). Continuous-time samplers leverage the fact that a Langevin SDE has the target density $\rho(x)$ as its stationary distribution, meaning that as $t\to\infty$, the law of the SDE converges to $\rho(x)$. However, in practice, one must discretize the SDE to simulate it, and such discretizations typically introduce bias and no longer preserve $\rho(x)$ exactly. To correct this bias, the discretized SDE can be used as a proposal within a Metropolis--Hastings framework, yielding a Markov chain that converges to the correct target. More broadly, the Metropolis--Hastings algorithm provides a general recipe for constructing discrete-time samplers that converge to $\rho(x)$. In high-dimensional settings, however, naive proposals can lead to very low acceptance rates, resulting in poor mixing and inefficient exploration of the target.
@@ 3.1 Discrete-Time Samplers
The classic Metropolis--Hastings algorithm [14, 30] constructs a Markov chain $\{X_{k}\}_{k\geq 0}$ with stationary distribution $\rho(x)$. Each iteration consists of two steps: (i) Given the current state $X_{k}$, propose a new state $Y_{k}\sim Q_{C}(\cdot\mid X_{k})$, where $Q_{C}(\cdot\mid X_{k})$ is a proposal distribution parameterized by some matrix $C$ (e.g., a covariance); (ii) Accept the proposal with probability
@@ 3.2 Continuous-Time Samplers
4 A Two-System Approach for McKean--Vlasov Equations
We now develop the two-system approach for general continuous-time mean-field SDEs (McKean--Vlasov equations), of which the overdamped and underdamped Langevin equations are special cases. The key idea is to consider two coupled systems of particles such that each system “drives” the evolution of the other. In a Langevin setting, this symmetric coupling is constructed to preserve the target distribution and to directly connect the behavior of the finite-$N$ particle system with its $N\to\infty$ mean-field limit.
The mean-field Langevin equation (11) is an example of a more general McKean--Vlasov SDE, which takes the form:
@@ 4 A Two-System Approach for McKean--Vlasov Equations
with $W_{t}^{1},\ldots,W_{t}^{N}$ independent Brownian motions and $\mathbf{X}_{t}=\{X_{t}^{j}\}_{j=1}^{N}$. As $N\to\infty$, this interacting particle system converges to the mean-field equation (14) in the sense of propagation of chaos [9, 42]: the particles become asymptotically independent and each follows the law of the McKean--Vlasov solution $\overline{X}_{t}$.
@@
Under Assumptions 1 and 2, one can show that the coupled SDE (18) has a unique strong solution on $[0,T]$ (proof in Appendix A.2). Moreover, $\textrm{Law}(Z_{1,t})$ and $\textrm{Law}(Z_{2,t})$ both coincide with the original mean-field law $\mu_{t}$ for all $t\in[0,T]$. In particular, if $Z_{1,0}$ and $Z_{2,0}$ are identically distributed (with law $\mu_{0}$), then $Z_{1,t}\stackrel{{\scriptstyle d}}{{=}}Z_{2,t}$ for all $t\in[0,T]$, and this common distribution evolves according to the McKean--Vlasov equation (14) (proof in Appendix A.3). Thus, $\pi_{t}=\textrm{Law}(Z_{t})$ can be seen as the coupling of two independent copies of the mean-field solution.
@@ 5 A Two-System Approach for Samplers
For the rest of this paper, unless otherwise stated, we consider a restricted class of McKean--Vlasov equations
$d\overline{X}_{t}\ =\ C(\mu_{t})b(\overline{X}_{t})\,dt\ +\ \sqrt{2C(\mu_{t})}\,dW_{t},\qquad t\in[0,T],$ (22)
where $\mu_{t}=\textrm{Law}(\overline{X}_{t})$, and $C(\mu)$ is Lipschitz continuous, bounded, and positive definite for all $\mu\in\mathcal{P}(\mathbb{R}^{d})$. In particular, $b(\overline{X}_{t})$ no longer depends on $t$ and the law of $\overline{X}_{t}$. As such, for $b(x)=-\nabla V(x)$, the invariant distribution takes the form $\rho(x)\propto\exp(-V(x))$. Thus, the two-system framework introduces a pair of interacting Markov chains, where each chain generates proposals based on the empirical law of the other. When applied to a continuous-time mean-field sampler, this coupling preserves $\rho$ as the invariant distribution in both the finite-particle setting and the infinite-particle limit. By adding Metropolis--Hastings acceptance steps to an alternating (staggered) time-discretization of this coupled system, we obtain an ensemble-chain sampler that exactly maintains the target distribution $\rho$.
This connection demonstrates that ensemble-chain samplers approximate expectations with respect to $\rho$ by leveraging particle interactions. In the limit as $N\to\infty$ (and as time $t\to\infty$), the ensemble statistics converge to their mean-field counterparts. This also motivates the use of adaptive samplers, which replace the whole ensemble with a single chain and rely on long-run time averages to recover the same mean-field behavior.
@@ 5.1 Continuous-Time Samplers
To illustrate, consider the mean-field overdamped Langevin example (11). The two-system construction leads to a coupled SDE on $Z_{t}=\{X_{t},Y_{t}\}$ (with $X_{t},Y_{t}\in\mathbb{R}^{d}$):
@@
for $t\in[0,T]$, where $\mu_{t}=\textrm{Law}(Z_{t})$, and $\rho(X_{t})$
Under mild regularity conditions on $\nabla\log\rho$, the two-system McKean--Vlasov equation (23) admits a unique strong solution (e.g., if $\log\rho(x)$ is strictly concave outside a large ball).
@@
Then $\nabla\cdot S(Z_{t})=0$ is a sufficient condition for the corresponding Fokker–Planck to have $\rho^{\otimes 2N}$ as the invariant density [32, 34]. Since $S(Z_{t})_{i,j}$ does not depend on $(Z_{t})_{j}$, $(\nabla\cdot S(Z_{t}))_{i}=0$ for all $i\in\{1,\ldots,2N\}$. Thus $\rho^{\otimes 2N}$ is the invariant density of (24). ∎
@@
Remark 2.
Lemma 1 can be extended to the case where the McKean–Vlasov equation takes the form
$d\overline{X}_{t}\ =\ C(\overline{\mu}_{t})b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sqrt{2C(\overline{\mu}_{t})}\,\sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t},$ (27)
and $\overline{\mu}_{t}=\textrm{Law}(\overline{X}_{t})$. In this case, the coupled SDE’s look like
$d\overline{X}_{t}\$$=\ C(\overline{\rho}_{t})b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sqrt{2C(\overline{\rho}_{t})}\,\sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t},$ (28)
$d\overline{Y_{t}}\$$=\ C(\overline{\mu}_{t})b(t,\overline{Y}_{t},\overline{\rho}_{t})\,dt\ +\ \sqrt{2C(\overline{\mu}_{t})}\,\sigma(t,\overline{Y}_{t},\overline{\rho}_{t})\,dW_{t},$ (29)
where $\overline{\rho}_{t}=\textrm{Law}(\overline{Y}_{t})$. As such, $\overline{Y}_{t}$ is only used to precondition the dynamics of $\overline{X}_{t}$, but is not used to compute $b$ or $\sigma$. As in Lemma 1, this satisfies $\nabla\cdot S(Z_{t})=0$, thus preserving the invariant density $\rho^{\otimes 2}$.
@@ 5.2 Discrete-Time Samplers
where $\mu_{t}^{Y}$ denotes the empirical law of subsystem $Y$ at time $t$, and $\xi^{X},\xi^{Y}$ are independent standard normal increments. In other words, we first update $X$ using the current $Y$-statistics, then update $Y$ using the new $X$-statistics (this is a symmetric split step akin to a leapfrog). This yields a sequence of proposals $(X_{t},Y_{t})\to(X_{t+h},Y_{t})\to(X_{t+h},Y_{t+h})$ for the coupled system. We then apply a Metropolis--Hastings accept/reject step at each proposal to ensure detailed balance with respect to $\rho$.
@@ 6 Algorithmic Realizations of Two-System Samplers
For the ensemble-chain (multi-chain) setting, convergence to the target $\rho$ is ensured as long as each Metropolis--Hastings update satisfies detailed balance. We therefore present the two-system ensemble versions of MALA and MAKLA in Algorithms 1 and 2.
@@ 7.1 Synthetic Experiments
Coupled MAKLA delivers order-of-magnitude gains over NUTS by exploiting spatial averaging, and the adaptive variants achieve similar per-gradient efficiency, often within $10\text{–}25$% of coupled MAKLA’s ESS/Grad, while using far fewer particles ($N=20$ chains versus $8d$ particles for Coupled). Under the two-system lens, this is natural: adaptive MAKLA replaces particle averaging with historical averaging. Instead of computing the preconditioner from a large, simultaneous ensemble, it estimates the same quantity from the chain’s time history. Thus, coupled and adaptive are two ways to approximate the same mean-field preconditioner; one by averaging across particles at a fixed time, the other by averaging across time with many fewer particles. Practically, this trades memory and ensemble size for a lightweight adaptation schedule, retaining most of the efficiency benefits of the coupled ensemble while substantially reducing particle count.
@@ 9 Acknowledgements
We thank Dr. Anna Lisa Varri, Dr. Amanda Lenzi, Prof. Peter Radchenko, Dr. Gonçalo dos Reis, and Dr. Deven Sethi for the helpful conversations while writing this paper. James Chok acknowledges the use of the Isaac Newton Trust (INT) grant G101121 LEAG/929. Daniel Paulin’s research was supported by an NTU Start Up Grant.
@@ A.2 Existence and uniqueness
This proof follows the standard argument of the existence and uniqueness of McKean--Vlasov SDEs, see [7, 21, 42] for more details.
@@
$\lvert\mathbf{b}(t,Z_{t},\pi_{t})-\mathbf{b}(t,\overline{Z}_{t},\overline{\pi}_{t})\rvert^{2}\$$\leq\ 2c\Big(\lvert Z_{t}-\overline{Z}_{t}\rvert^{2}\ +\ \mathcal{W}_{2}^{2}(p_{1}\circ\pi_{t},p_{1}\circ\overline{\pi}_{t})+\mathcal{W}_{2}^{2}(p_{2}\circ\pi_{t},p_{2}\circ\overline{\pi}_{t})\Big)$ (51)
@@ A.3 Convergence in Wasserstein
If $\textrm{Law}(X_{0})=\textrm{Law}(Y_{0})=\textrm{Law}(\overline{X}_{0})=\textrm{Law}(\overline{Y}_{0})$, then $\textrm{Law}(X_{t})=\textrm{Law}(Y_{t})=\textrm{Law}(\overline{X}_{t})=\textrm{Law}(\overline{Y}_{t})$ for $t\in[0,T]$. That is, the mean-field dynamics of the two-system approach are the same as using one particle.
@@
$\begin{split}&dX_{t}\ =\ b(t,X_{t},\rho_{t})\,dt\ +\ \sigma(t,X_{t},\rho_{t})\,dW_{t}^{X},\\ &dY_{t}\ =\ b(t,Y_{t},\mu_{t})\,dt\ +\ \sigma(t,Y_{t},\mu_{t})\,dW_{t}^{Y},\\ &d\overline{X}_{t}\ =\ b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t}^{X}\\ &d\overline{Y}_{t}\ =\ b(t,\overline{Y}_{t},\overline{\rho}_{t})\,dt\ +\ \sigma(t,\overline{Y}_{t},\overline{\rho}_{t})\,dW_{t}^{Y}\end{split}$ (70)
with $\textrm{Law}(X_{t})=\mu_{t},\,\textrm{Law}(Y_{t})=\rho_{t},\,\textrm{Law}(\overline{X}_{t})=\overline{\mu}_{t},\,\textrm{Law}(\overline{Y}_{t})=\overline{\rho}_{t}$ and initial conditions $X_{0},Y_{0},\overline{X}_{0},\overline{Y}_{0}$ respectively. Then for $t\in[0,T]$,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\Big[\mathcal{W}_{2}^{2}(\mu_{0},\overline{\mu}_{0})\ +\ \mathcal{W}_{2}^{2}(\rho_{0},\overline{\rho}_{0})\ +\ \mathcal{W}_{2}^{2}(\overline{\mu}_{0},\overline{\rho}_{0})\Big].$ (71)
@@
$\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,X_{s},\rho_{s})-\sigma(s,\overline{X}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (75)
@@
Since a similar inequality holds for $\mathbb{E}\lvert Y_{t}-\overline{Y}_{t}\rvert^{2}$,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})\$$\leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{Y}_{0}\rvert^{2}\right)$ (79)
$\qquad\qquad+c\int^{t}_{0}\Big(\mathcal{W}_{2}^{2}(\rho_{s},\overline{\mu}_{s})\ +\ \mathcal{W}_{2}^{2}(\mu_{s},\overline{\rho}_{s})\Big)ds.$ (80)
$\leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{Y}_{0}\rvert^{2}\right)$ (81)
$\qquad\qquad+c\int^{t}_{0}\Big(\mathcal{W}_{2}^{2}(\rho_{s},\overline{\rho}_{s})\ +\ \mathcal{W}_{2}^{2}(\mu_{s},\overline{\mu}_{s})\ +\ \mathcal{W}_{2}^{2}(\overline{\rho}_{s},\overline{\mu}_{s})\Big)ds,$ (82)
using the triangle inequality. Since $(X_{t})_{0\leq t\leq T}$, $(Y_{t})_{0\leq t\leq T}$, $(\overline{X}_{t})_{0\leq t\leq T}$, and $(\overline{Y}_{t})_{0\leq t\leq T}$ are squared-integrable, it follows that $\mathbb{E}\lvert X_{t}-\overline{X}_{t}\rvert^{2}$ and $\mathbb{E}\lvert Y_{t}-\overline{Y}_{t}\rvert^{2}$ are bounded for $t\in[0,T]$. Hence $\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})+\mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})+\mathcal{W}_{2}^{2}(\overline{\mu}_{t},\overline{\rho}_{t})$ is also bounded on $t\in[0,T]$. This allows us to apply Grönwall’s inequality
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})\ \leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{X}_{0}\rvert^{2}\ +\ \int^{t}_{0}\mathcal{W}^{2}_{2}(\overline{\rho}_{s},\overline{\mu}_{s})\,ds\right).$ (83)
A similar argument can be used to show that $\mathcal{W}^{2}_{2}(\overline{\mu}_{t},\overline{\rho}_{t})\leq c\mathbb{E}|\overline{X}_{0}-\overline{Y}_{0}|^{2}$. Thus,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})\ \leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert\overline{Y}_{0}-\overline{X}_{0}\rvert^{2}\right).$ (84)
@@
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\Big[\mathcal{W}_{2}^{2}(\mu_{0},\overline{\mu}_{0})\ +\ \mathcal{W}_{2}^{2}(\rho_{0},\overline{\rho}_{0})\ +\ \mathcal{W}_{2}^{2}(\overline{\mu}_{0},\overline{\rho}_{0})\Big].$ (85)